help me
1 Solve the egua rim : y - 2y + 2y 0• (20) 2 Solve the equotion: y - 4y + 5y = 0 y(0) = 2 y(0)= 6 • (20) 3 Solve the lin
Answer #1

3. x + 2x2+1 234 + 5x2 + x + 0x2 - Co-efficie- Ari 2 y-2y+zy ... - Let yo ex be the trial solub of Then auxiliary egun of isX + 2X2 +373 7. 224 + 5x2 + 3x3 = 12 x + 0x2 + 8x3 =10 Damat . Coefficient matrix of sy Ari 237 tem of egun ② is 12 5 3 II 0. Ja is date the seatstenot Hence the solue of system of eque 34= 2X2 = 1, 43=17 when wena A Let A be the eigen valor of . [AThe corresponding eigen vector (2-2) is When x = 3. The corresponding eigen sector as lh we have F4-3 -1747 E 1-3 1 2] → Y-42